OpenStudy (anonymous):

Solve the System. Give your answer as (X,Y,Z) -5x-4y-4z=9 -6x+5y-2z=-15 -2x-3y-2z=13 Please i need help

OpenStudy (anonymous):

subtract the third equation from the second, which will give \[-4x+8y=-28\]

OpenStudy (anonymous):

then multiply the third equation by \(-2\) and add to the first equation you get \[-5x-4y-4z=9\]\[4x+6y+4z=-23\] add and get \[-x+2y=14\]

OpenStudy (anonymous):

now you can solve the system \[-x+2y=14\]\[-4x+8y=28\]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

yea i got stuck there i divide the bottom by -4

OpenStudy (anonymous):

yeah i guess this does not have a solution

OpenStudy (anonymous):

- 5x _ 4y - 4z = 9 -->(5)-5x - 4y - 4z = 9 -6x + 5y - 2z = - 15 ->(4)-6x + 5y - 2z = - 15 -------------------- -25x - 20y - 20z = 45 (result of multiplying by 5) - 24x + 20y - 8z = - 60 (result of multiplying by 4) --------------------add - 49x - 28z = - 15 -6x + 5y - 2z = - 15 -->(3)-6x + 5y - 2z = - 15 -2x - 3y - 2z = 13--> (5)-2x - 3y - 2z = 13 ---------------------- -18x + 15y - 6z = - 45 (result of multiplying by 3) -10x - 15y - 10z = 65 ----------------------add - 28z - 16z = - 20 - 28x - 16z = - 20 -->(7)-28x - 16z = - 20 -49x - 28z = - 15 -->(-4)-49x - 28z = - 15 --------------------- - 196x - 112z = - 140 (result of multiplying by 7) 196x + 112z = 60 ---------------------add 0 + 0 = - 80 your right satellite73......no solution :)